OpenStudy (anonymous):

PLEASE HELP ME I WILL FAN AND MEDAL!!! The equation y = 13x shows that a boat travels at a speed of 13 km/h. Which ordered pair satisfies the equation y = 13x? A. (13, 0) B. (13, 1) C. (1, 14) D. (3, 39)

OpenStudy (anonymous):

@Nnesha @welshfella

OpenStudy (welshfella):

just plug in the values of x and y and see if they fit

OpenStudy (welshfella):

(13,0) means x = 13 and y = 0 so see if this one fits y = 13x 0 = 13*13 = 169 No way!

OpenStudy (welshfella):

so try B C and D

OpenStudy (welshfella):

B replace the x by 13 and the y by 1 y = 13x y = 1 x = 13 so 1 = 13* 13 = 169 so?

OpenStudy (welshfella):

do you understand what i'm doing the value for y must equal the value for 13 x

OpenStudy (welshfella):

for A and B they do not

OpenStudy (anonymous):

@Nnesha

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi

OpenStudy (misty1212):

pick the one where the second coordinate is 13 times the first coordinate

OpenStudy (misty1212):

you need another hint?

OpenStudy (anonymous):

D.?

OpenStudy (misty1212):

Taylor Swift does not need to do algebra!

OpenStudy (anonymous):

Lol

OpenStudy (misty1212):

\[3\times 13=39\]pick that one

OpenStudy (misty1212):

yeah D

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

@misty1212 how do you make that heart?

OpenStudy (misty1212):

my secret!

OpenStudy (misty1212):

actually, if you want to see any code, math or otherwise, right click on it and select "show math as" \(\to\) "latex"

OpenStudy (anonymous):

Ok

OpenStudy (misty1212):

actually "tex commands" not latex